A Clock Face has negative charges-q,-2q,-3q......-12q,fixed at the positionof the corresponding numeral of the dial the clock hand do not disturb the netfield due to point charges at what time does the hour hand point in the same direction is electric field at the centre of dial

Grade:12

2 Answers

Divyanshu Rawat
6 years ago
Here attractive force will workSo if we assume a test charge placed at centre then for numbers of right side means in opposite of -12q we have -6q so half of the force will be cancelled and the net due to both will be in direction of -12q charge.Similarly this will be followed by other number too.So the the resultant of all net forces will be somewhere b/w no. 9 and 10.So at 9:30 the hour hand will direct towards the direction of net Electric field.I hope this helps. Thanks a lot
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
When we measure the electric field at the center the opposite numbers each having a difference of 6 will have the same difference of charge, hence if we consider two opposite position as one pair. Each pair will give the same magnitude of electric field.
These resultant vectors of electric fields are directed towards
From 1 → 7
2 → 8
3 → 9
4 → 10
5 → 11
6 → 12
 
Now the resultant will lie in exactly the middle of all the angle.[as all inbetween angles as well as magnitude of field is equal]
Hence, the resultant will be directing towards, (12 + 7)/2 = 19/2 = 9.5
Now 9.5 on clock represents 9:30
 
Thanks and regards,
Kushagra